calculate the current through the galvanometer and obtain the effective resistance of the network


1
Expert's answer
2021-10-18T11:04:59-0400


resistance of galvanometer is "17\ \Omega"


For loop ABDA:

"55I_{AB}+17I_G-21I_{BD}=0"


For loop BCDB:

"25(I_{AB}-I_G)-45(I_{BD}+I_G)-17I_G=0"


For loop ADCEA:

"21I_{BD}+45(I_{BD}+I_G)=24"


So, we have:

"55I_{AB}+17I_G-21I_{BD}=0"

"25I_{AB}-87I_G-45I_{BD}=0"

"66I_{BD}+45I_G=24\implies 22I_{BD}+15I_G=8"


"I_{BD}=0.36-0.68I_G"


"55I_{AB}+31.28I_G=7.56"

"25I_{AB}-56.4I_G=16.2"


"I_{AB}=0.65+2.26I_G"

"35.75+124.3I_G+31.28I_G=7.56"


the current through the galvanometer:

"I_G=(7.56-35.75)/(124.3+31.28)=-0.18\"A


the effective resistance of the network:

"\frac{1}{R}=\frac{1}{17}+\frac{1}{45+25}+\frac{1}{55+21}=\frac{}{}=\frac{5320+1292+1190}{90440}=\frac{7802}{90440}"


"R=\frac{90440}{7802}=11.6\ \Omega"